Two points X and Y, 1000 m apart are located on a straight road running east and west. From



X, the bearing of a tower is 32°W of N and from Y the bearing of the town 26° N of E. Determine



the shortest distance of the tower from the road.

1
Expert's answer
2021-11-11T18:40:37-0500

i....Cos32°=100/d

d=100/Cos32°

d=117.9m

ii.....Cos26°=100/d

d=100/Cos26°

d=111.3m

Therefore, the shortest distance = 111.3m